# Data-Efficient Sleep Staging with Synthetic Time Series Pretraining

### v1

@@ Introduction
Figure 1: The training process consists of a pretraining and a fine-tuning phase. In the pretraining phase, the feature extractor $f$ produces features $h_{i}$ and is trained together with a classifier $c_{p}$ to detect the frequency content of randomly generated synthetic time series signals (multi-label classification problem). In the fine-tuning phase, the pretrained feature extractor $f$ extracts features $h_{i}$ from individual epochs of EEG and EOG signals. The features of a sequence of epochs (training sample) are then aggregated by a classifier $c_{f}$ to predict the sleep stage of the middle epoch in the sequence (multi-class classification problem).
Methods
Fig. 1 presents an overview of our training scheme, which comprised two phases. In the pretraining phase, we generated synthetic time series signals and trained a convolutional feature extractor with a multi-layer perceptron classifier to predict the frequency content of these signals. In the fine-tuning phase, we utilized the pretrained feature extractor together with another classifier to perform sleep staging on EEG and EOG (electrooculography) signals.
Data
Synthetic Data
For the pretraining phase (see Fig. 1), we defined a simple random process to generate synthetic time series signals. Each synthetic signal $s_{c}(t)$ was a normalized time series composed of the sum of 30-second sine waves sampled at 100 Hz with random frequencies and phases,
$s_{c}(t)=\frac{\tilde{s}_{c}(t)-\mu_{c}}{\sigma_{c}},\quad\tilde{s}_{c}(t)=\sum_{i}^{n_{f}}I_{i}\,\sin(2\pi f_{c,i}+\phi_{i}),\quad I_{i}\in\mathcal{U}\{0,1\}$ (1)
where $t$ denotes time (in seconds), $n_{f}$ denotes the number of frequency bins, $I_{i}$ denote whether a frequency bin is used (1) or not (0), $f_{c,i}$ denote the frequencies of the sine waves (in Hertz), $\phi_{i}$ denote their phase (in radians), and $\mu_{c}$ and $\sigma_{c}$ denote mean and standard deviation of $\tilde{s}_{c}$. The phases $\phi_{i}$ of the sine waves were sampled from a uniform distribution between 0 and $2\pi$. To sample the frequencies $f_{c,i}$, we first divided the frequency range of 0.3–35 Hz recommended by the American Academy of Sleep Medicine (AASM) for filtering EEG and EOG signals into $n_{f}=20$ bins with a base 2 logarithmic scale (see x-axis of Fig. 4c for the resulting frequency bins). We chose a logarithmic scale for the frequency bins because typical power spectral densities of EEG and EOG signals show more power in lower frequency bands than in higher bands. We then randomly decided for each frequency bin (with a probability of 50% for each bin) whether it would be used to create the synthetic signal ($I_{i}=1$) or not ($I_{i}=0$). Within each selected frequency bin, we randomly sampled the final frequencies $f_{c,i}$ of the sine waves. Finally, the generated sine waves were summed and normalized to create synthetic signals $s_{c}(t)$.
When pretraining our neural networks, each training sample $s(t)=(s_{1}(t),s_{2}(t),s_{3}(t))$ consisted of three synthetic signals, corresponding to three “channels” of sleep staging data, and an associated label vector $l=(I_{0},\dots,I_{n_{f}})$. The three synthetic signals of a sample shared the same phases and frequency bins. However, the frequencies of the sine waves were sampled independently for each channel. The label vector $l$ encoded the frequency bins from which the frequencies of the sine waves were drawn in a one-hot encoded format. Pretraining involved predicting all frequency bins encoded in this label vector $l$, which made it a multi-label classification problem with $n_{f}=20$ classes. For each training run, we generated 101,000 synthetic samples, using 100,000 samples for training and the remaining 1000 samples for tracking various metrics during training. The number of synthetic samples, frequency bins, and the logarithmic scale that defines the boundaries between the bins were determined through preliminary experiments where we explored different hyperparameters.
Sleep Staging Data
During the fine-tuning phase (see Fig. 1), we used two publicly available datasets: the DODO and DODH datasets [20]. The DODO dataset contains 55 recordings from subjects diagnosed with obstructive sleep apnea (OSA), while the DODH dataset contains 25 recordings from healthy subjects. Each recording stems from a different subject and was split into 30-second non-overlapping windows (epochs), which were annotated with sleep stages (Wake, N1, N2, N3, REM) by five sleep experts following the AASM guidelines [17]. Inspired by the work of Guillot et al. [20], we used the annotations of all five scorers to create a consensus annotation for each epoch. To create this consensus, we selected the most voted stage for each epoch. In cases of ties, we chose the sleep stage determined by the most reliable scorer. For each recording, the most reliable scorer was identified as the scorer with the highest average agreement with all other scorers. In our experiments, we focused on the following EEG and EOG derivations that are available in both datasets and follow the AASM recommendations [17]: C3_M2, F3_M2, EOG1. All channels were sampled at 250 Hz.
To prepare the data for training, we filtered the signals between 0.3 and 35 Hz with an 8th-order Butterworth filter and downsampled the filtered signals to 100 Hz using polyphase filtering. Next, we removed all epochs from the data that were annotated as artifacts by the consensus of the sleep experts. We then normalized the amplitudes of each individual epoch by subtracting the mean and dividing the result by the interquartile range of its amplitude distribution. Finally, we clipped all signal amplitudes above 20 or below -20 to minimize outliers.
After preprocessing the signals, we combined the DODO and DODH datasets into a single dataset of 80 recordings. We then split the data subject-wise into 5 folds for cross-validation using stratified sampling. Consequently, each fold contained 16 recordings, with 11 recordings from the DODO dataset and 5 recordings from the DODH dataset. In addition to splitting the data into training and test folds, we reserved two recordings from each fold as validation data for hyperparameter tuning and early stopping. We validated our models on the validation data of the four training folds. The validation data of the test fold was kept separate and not involved in validation or testing.
Model

### v2

@@ 2.1 Training Configurations
Figure 1: The training process consists of a pretraining and a fine-tuning phase. In the pretraining phase (“frequency pretraining”), we generated synthetic time series signals by summing sine waves with random frequencies. These synthetic signals were used to train a deep neural network consisting of a feature extractor $f$ and a classifier $c_{p}$ to predict the frequencies present in the signals (multi-label classification problem). In the fine-tuning phase, the pretrained feature extractor $f$ produced features $h_{i}$ from individual EEG and EOG epochs. The features of a sequence of these epochs were then used by a classifier $c_{f}$ that was trained to predict the sleep stage of the middle epoch in the sequence (multi-class classification problem).
Fully Supervised. The Fully Supervised training configuration is similar to many existing deep learning approaches for sleep staging [36] and served as a baseline to compare our pretrained models against. In this configuration, we skipped the pretraining step and trained (fine-tuned) the feature extractor $f$ and classifier $c_{f}$ from scratch using sleep staging data.
Fixed Feature Extractor. We employed the Fixed Feature Extractor configuration to investigate the relevance of the features generated by the pretrained feature extractor for sleep staging. After pretraining the feature extractor $f$ on synthetic data, we kept its model weights and BatchNorm statistics fixed and only fine-tuned the sleep staging classifier $c_{f}$ on sleep data.
Fine-Tuned Feature Extractor. With this training configuration, we studied (i) how model performance changes when the pretrained feature extractor is allowed to change during fine-tuning and (ii) whether the priors learned during pretraining can prevent overfitting in few-sample or few-subject regimes. As in the previous configuration, we first pretrained the feature extractor $f$ on synthetic data, but then fine-tuned both the feature extractor and the classifier $c_{f}$ on sleep data without keeping any model weights fixed. Consequently, this configuration is similar to the Fully Supervised configuration with the key distinction that the feature extractor is initialized with pretrained weights.
Untrained Feature Extractor. The Untrained Feature Extractor configuration served as a baseline to study whether our pretraining scheme produces priors that are superior to random weights for sleep staging. We randomly initialized the feature extractor $f$ using Kaiming normal initialization [24] and then kept its weights fixed while fine-tuning the classifier $c_{f}$. This approach mirrors the Fixed Feature Extractor configuration, but with a random feature extractor instead of a pretrained one.
2.2 Data Efficiency
To assess the data efficiency of our pretraining method, we compare the performance of the different training configurations when fine-tuned with a reduced amount of training data (low-data regime) or the full training data (high-data regime). We investigated the low-data regime by first pretraining models with the full synthetic dataset. Depending on the training configuration, we then fine-tuned the pretrained or randomly initialized models with the data of 50 randomly sampled sleep staging samples from one subject. The sampling procedure selected sleep staging data without class stratification, and each sleep stage was represented at least once in the reduced datasets. In the high-data regime, we followed the same procedure but fine-tuned models with the full training data instead of a reduced amount of data. We repeated all experiments three times in a 5-fold cross-validation scheme, resulting in 15 training runs for each configuration and dataset. This approach allowed us to estimate the spread of the macro F1 scores when using different model initializations, sleep staging samples, and data folds for training and testing (see Section 4).
In the low-data regime, we observed that models pretrained with synthetic data outperformed models trained from scratch in sleep stage classification (see Figure 2). The performance gap between pretrained and non-pretrained models was most pronounced when comparing the fine-tuned feature extractor to the fully supervised configuration, with the former achieving average macro F1 scores that were 0.06–0.07 higher than those of the latter across datasets. When removing the fine-tuning of the pretrained features (fixed feature extractor configuration), our method still yielded macro F1 scores that were 0.01–0.05 higher than those of the fully supervised configuration. Comparing the fixed feature extractor to the untrained feature extractor configuration further highlights the importance of the learned features. Pretraining the feature extractor with synthetic data improved the macro F1 scores by 0.10–0.17 compared to a random initialization. While the macro F1 scores of the training configurations varied between datasets, the general trends observed in the low-data regime were consistent across all three datasets.
In the high-data regime, our pretrained models were on par with fully supervised trained models (p ¡ 0.05 for a paired TOST test with a margin of $\pm 0.01$ on the DODO/H and ISRUC datasets) and achieved competitive performance in sleep stage classification (see Figure 2). The fine-tuned feature extractor configuration achieved average macro F1 scores of 0.76–0.81 across datasets, comparable to the macro F1 scores of 0.76–0.80 achieved by the fully supervised configuration. As in the low-data regime, we observed that fine-tuning the pretrained features was beneficial, as the macro F1 scores achieved by the fine-tuned feature extractor configuration were 0.02–0.05 higher than those of the fixed feature extractor configuration. The performance gap between the untrained feature extractor and the fixed feature extractor remained substantial even in the high-data regime, with the former achieving average macro F1 scores that were 0.08–0.12 lower than those of the latter.
Figure 2: Average macro F1 scores achieved by the different training configurations when fine-tuned with a small set of training data (left side) or a large set (full set) of training data (right side) from one of the three datasets DODO/H, Sleep-EDFx, and ISRUC. The bars indicate the mean of the macro F1 scores measured on the test sets and averaged over 15 trainings (3 repetitions of a 5-fold cross-validation). Error bars show the standard deviation of the macro F1 scores, and * symbols indicate significant differences between Untrained and Fixed Feature Extractor or between Fine-Tuned Feature Extractor and Fully-Supervised training configuration $(p<0.01)$ according to a one-sided Wilcoxon signed-rank test.
2.3 Impact of Subject Diversity and Number of Training Samples
We further explored the data efficiency of our pretraining method by investigating how model performance is affected by subject diversity and sample volume (i.e., number of samples) in the training data used for fine-tuning. To study the effect of subject diversity independent of sample volume, we separately varied the number of subjects and the number of samples in the training data. The number of subjects was randomly sampled as $n_{subj}\in\{1,2,3,4,5\}$, and the number of samples was randomly sampled from those subjects as $n_{samples}\in\{50,130,340,900,\text{all}\}$ (“all” indicates that all available training samples were used). Our sampling strategy selected sleep staging data without class stratification, and each sleep stage was represented at least once in the reduced datasets. For each parameter combination, we trained the fully supervised and fine-tuned feature extractor configurations on all three datasets with three repetitions of 5-fold cross-validation.
In our results, we observed that both the fully supervised and the fine-tuned feature extractor configurations benefited from increased subject diversity, even when the total number of training samples was held constant (see rows in Figure 3a–f). Similarly, both configurations benefited from an increased number of training samples when the number of subjects was held constant (see columns in Figure 3a–f). For the considered parameter ranges, the impact of reduced subject diversity and reduced sample volume on model performance appeared to be comparable.
Figure 3: Detailed comparison of the fully supervised and fine-tuned feature extractor configurations in few-sample and few-subject regimes for each of the three datasets DODO/H (panels a,d,g), Sleep-EDFx (matrices b,e,h), and ISRUC (panels c,f,i). Panels (ac) and Panels (df) show the macro F1 scores achieved by the fully supervised and the fine-tuned feature extractor configurations, respectively. The displayed scores are the mean and the standard deviation of the macro F1 scores measured on the test set over 15 trainings (3 repetitions of a 5-fold cross-validation). Panels (g–i) show the average macro F1 scores achieved by the fine-tuned feature extractor configuration minus the average macro F1 scores achieved by the fully supervised configuration. These differences were calculated using a bootstrapping approach with 10,000 bootstrap samples. In this bootstrapping approach, we first paired the 15 macro F1 scores available for both configurations for each matrix entry based on the used seed and fold. We then calculated the difference between each pair of macro F1 scores. For each bootstrap sample, we sampled 15 of these differences with replacement and calculated their average. Finally, we display the mean and the standard deviation of these bootstrap samples.
Similar to the observations made in Section 2.2, the fine-tuned feature extractor configuration achieved better macro F1 scores than the fully supervised configuration in the few-sample regime. The performance gap between the two configurations was most evident when the number of samples was limited to 50, with bootstrap estimates of the mean differences in macro F1 scores ranging from 0.05 to 0.09 across datasets and subject numbers (see Figure 3g–i). These differences in macro F1 scores decreased as the number of samples increased. When training on all available training samples, the fine-tuned feature extractor configuration achieving comparable or slightly better performance than the fully supervised configuration. Interestingly, the performance gap between the two configurations with all available training samples was most pronounced for the DODO/H dataset. The fine-tuned feature extractor configuration achieved average macro F1 scores that were 0.02–0.06 higher than those of the fully supervised configuration. For the Sleep-EDFx and ISRUC datasets, the performance differences between the two configurations with all available training samples were minimal at 0.00–0.01.
Depending on the dataset, the fine-tuned feature extractor configuration showed varying degrees of improvement over the fully supervised configuration in the few-subject regime. When training with data of a single subject, the fine-tuned feature extractor configuration achieved average macro F1 scores that were 0.06–0.11 higher across sample numbers than those of the fully supervised configuration for the DODO/H dataset (see rows in Figure 3g). This performance gap decreased as the number of subjects increased, with mean bootstrapped differences in macro F1 scores between 0.02 and 0.09 for five subjects. In contrast, varying the number of subjects had less impact on the performance gap between the fine-tuned feature extractor and the fully supervised configurations for the Sleep-EDFx and ISRUC datasets (see rows in Figure 3h–i). The improvements achieved by the fine-tuned feature extractor configuration when trained with one subject (0.01–0.05 for the Sleep-EDFx dataset and 0.00–0.07 for the ISRUC dataset) were comparable to those achieved when trained with five subjects (0.00–0.06 for the Sleep-EDFx dataset and 0.01–0.06 for the ISRUC dataset).
2.4 Priors Towards Frequency Information
To get a better understanding of the pretraining process and the priors learned by the model, we recorded several metrics during pretraining. The recorded loss function converged to a low value, indicating that the model has learned effectively (see Figure 4a). At the same time, the Hamming metric reached a high value of 0.9 on the validation data (see Figure 4b). This value can be interpreted as the model predicting the frequency bins that were used to create the synthetic signals with an accuracy of 90%. The model was particularly proficient in predicting higher frequencies starting from 2.5 Hz (accuracies $90\%$; see Figure 4c). Lower frequencies, especially those below 1 Hz, were predicted with lower accuracy (75–85%).
We hypothesize that the differences in prediction accuracy across frequency bins are not due to the pretrained model being unable to predict lower frequencies. Instead, we believe this discrepancy arises from the varying width of the frequency bins (see x-axis of Figure 4c) as the bins for lower frequencies were narrower than those for higher frequencies due to their logarithmic scaling. While we applied a logarithmic binning scheme to increase our model’s focus on the lower frequencies important for identifying slow wave sleep (N2 and N3 sleep), the narrow low frequency bands increase the difficulty of the pretraining task leading to decreased accuracies. Preliminary exploration of this trade-off between model focus and task difficulty showed that the current binning scheme slightly outperforms a linear binning scheme in the fixed feature extractor setup.
Interestingly, the ability of the pretrained feature extractor to extract useful features for sleep staging was strongly influenced by the diversity of the synthetic pretraining data (see Figure 4d). We investigated this influence of sample diversity by pretraining models with varying amounts of different synthetic samples $n_{synthetic}\in\{1,10,100,10^{3},10^{4},10^{5},10^{6}\}$. To isolate the effect of sample diversity from the effect that the number of training steps has on model performance, we kept the number of gradient updates per training epoch constant by under- or oversampling the synthetic data to 100,000 samples. The pretrained models were then trained for sleep staging using data from one random subject from the DODO/H dataset and evaluated on the validation data of each cross-validation fold. As in the previous experiments, we performed three repetitions of a 5-fold cross-validation scheme for each number of synthetic samples.
Figure 4: Analysis of the “frequency pretraining” task. Panel (a) shows the development of the loss function during a single pretraining run for both the training and validation data. The following two panels quantify the accuracy of a pretrained model to predict the frequencies of the synthetic signals. The hamming metric in panel (b) measures the overall accuracy summarized across all frequency bins, while panel (c) shows the accuracy for each frequency bin separately. Panel (d) illustrates how the diversity of the synthetic pretraining data affects the performance of the pretrained feature extractor (FE) after fine-tuning for sleep staging on a single subject from the DODO/H dataset. Each bar indicates the mean of the macro F1 scores on the validation data averaged over 3 repetitions of a 5-fold cross-validation, while the error bars show the standard deviation of the macro F1 scores.
Both the fixed feature extractor and the fine-tuned feature extractor configurations required a high level of diversity in the synthetic pretraining data to achieve good sleep staging performance (see Figure 4d). When pretrained with only one synthetic sample, the performance of the two training configurations with pretraining differed only slightly from the performance of the configurations without pretraining (untrained feature extractor and fully supervised configurations). As the number of synthetic samples was increased to more than 100, the performance of the pretrained models improved substantially until reaching a plateau at around 10,000 samples. We hypothesize that this plateau was reached because the model had learned all the relevant features from the pretraining task, and additional samples did not provide any further benefit. The simplicity of the pretraining task could also explain the negligible performance differences between the pretrained configurations with less than 100 synthetic samples and the training configurations without pretraining. For such few synthetic samples, the model may have memorized the synthetic data rather than learn general features useful for sleep staging.
2.5 Comparison to Self-Supervised Methods
We compared our frequency pretraining approach (FPT) with two popular self-supervised learning (SSL) methods, SimCLR [9] and VICReg [4], using the same model architecture as in FPT. The projection head (SimCLR) and expander (VICReg) followed their original designs [9, 4]. Different from the conceptually simple multi-label classification task in FPT, both SSL methods follow the objective of contracting representations of data-augmented “positive” samples while keeping representations of dissimilar “negative” samples far apart (see original works for details [9, 4]). Positive samples were generated using standard data augmentations: amplitude scaling (random factor 0.5–2), Gaussian noise injection ($\sigma$ = 0.05), random temporal masking (10 segments of 1.5–3 s), time shifting (up to $\pm$1.5 s), and time warping (random factor 0.67–1.5). We evaluated two pretraining configurations: pretraining on (i) synthetic data (as in FPT) and (ii) EEG recordings from the ISRUC and Sleep-EDFx datasets. Fine-tuning was performed on DODO/H datasets under low-data (50 random epochs from one random subject) and high-data (all available training data) regimes, respectively.
When pretrained on synthetic data, both SimCLR and VICReg achieved performance comparable to the FPT method, with FPT slightly outperforming the SSL methods in the high-data regime when the feature extractor was fine-tuned (average MF1 scores of 0.80 versus 0.81, see Table 1). Pretraining on EEG data yielded only modest improvements, with both SSL methods showing slightly better performance in the low-data regime when the feature extractor was fixed (difference in average MF1 scores of 0.02) and in the high-data regime when fine-tuning the feature extractor (difference in average MF1 scores of 0.01). Compared to the EEG-pretrained SimCLR and VICReg models in the high-data regime, FPT showed similar performance (average MF1 scores of 0.76 and 0.81) but, unlike the two SSL methods, did not require EEG data for pretraining. In the low-data regime, the SSL methods performed comparable to the FPT method when fine-tuning the feature extractor and achieved slightly higher scores than FPT when keeping the feature extractor fixed (average MF1 scores of 0.44 and 0.45 versus 0.42). None of the differences in MF1 scores between the SSL methods and FPT were significant according to a two-sided Wilcoxon signed-rank test (p-values >> 0.05).
Table 1: Comparison of different pretraining methods when fine-tuned with 50 random sleep epochs of one subject (low-data regime) or the full training set (high-data regime) of the DODO/H dataset. During fine-tuning, the feature extractor was either kept fixed (FE fixed) or updated together with the classifier (FE fine-tuned). We pretrained SimCLR and VICReg on the same synthetic data used for the frequency pretraining task, in addition to the traditional approach of pretraining on EEG data (i.e., ISRUC and Sleep-EDFx datasets). MF1 scores were calculated as the average of three repetitions of a five-fold cross validation (standard deviation in parentheses).
Pretraining Method Avg. MF1 Low-Data Regime Avg. MF1 High-Data Regime
FE Fixed FE Fine-Tuned FE Fixed FE Fine-Tuned
SimCLR, pretrained with synth. data 0.42 (0.08) 0.44 (0.09) 0.76 (0.02) 0.80 (0.02)
SimCLR, pretrained with EEG data 0.44 (0.08) 0.43 (0.09) 0.76 (0.02) 0.81 (0.02)
VICReg, pretrained with synth. data 0.43 (0.07) 0.44 (0.08) 0.76 (0.02) 0.80 (0.02)
VICReg, pretrained with EEG data 0.45 (0.08) 0.45 (0.09) 0.76 (0.02) 0.81 (0.01)
Frequency pretraining 0.42 (0.06) 0.44 (0.09) 0.76 (0.01) 0.81 (0.02)
3 Discussion
In this work, we propose a novel pretraining scheme for EEG time series data that leverages synthetic data generated by a simple random process. We specialized the hyperparameters of the pretraining task to the typical frequency range and distribution of sleep EEG signals and demonstrated the effectiveness of this task for sleep stage classification. Due to the availability of several open sleep staging datasets [35, 47], we were able to fully control the amount and diversity of the training data, which allowed us to study the impact of our method in different data regimes. We hypothesize that our pretraining scheme could be particularly beneficial in few-sample and few-subject regimes, which we argue could benefit greatly from the priors towards frequency information that a model learns during pretraining.
Our results confirm the effectiveness of our pretraining scheme, particularly in few-sample and few-subject regimes. Pretrained models outperformed non-pretrained models when fine-tuned with a reduced number of subjects or training samples (see rows and columns in Figure 3g–i, respectively). The performance gap between pretrained and non-pretrained models was most pronounced in the few-sample regime, where pretrained models consistently achieved improvements over non-pretrained models across multiple datasets. In the few-subject regime, this performance gap was not as consistent across datasets, with our pretraining method showing the most substantial improvements in the DODO/H dataset. Our findings support observations made in the field of Self-Supervised Learning (SSL) that pretrained models generally have better data efficiency than fully supervised ones [2, 13]. In contrast to SSL methods, however, our pretraining scheme improves data efficiency without requiring empirical data, while achieving comparable or only slightly reduced performance (see Table 1). Interestingly, we observed that two data-augmentation-based SSL methods performed well even when pretrained with synthetic data instead of EEG data (see Table 1), suggesting that SSL approaches are promising, yet computationally intensive, alternatives to the FPT method. Further exploration of the use of SSL methods for pretraining on synthetic data is warranted and seems promising. Generating synthetic data can be much cheaper and more cost-effective than collecting EEG data, as generating 10,000 samples of the synthetic data used in this study required only $8.17\pm 0.37$ s of a single consumer-grade CPU core (Lenovo Legion S7 16IAH7 Laptop with an Intel® Core i7-12700H CPU).
We hypothesize that the potential of synthetic data stems from the priors that the model learns during pretraining. These priors could prevent overfitting to a small number of training samples, particularly those from minority classes (e.g., N1 sleep), or subject-specific features, which is especially problematic in situations with very little training data. As expected, we observed that all of our training configurations improved with a larger training dataset (see Figure 2). This aligns with the prevalent view in the literature that deep learning models for sleep staging need substantial amounts of diverse data to perform well [36, 1, 15, 14]. When trained with the full training data, pretrained models performed comparably to fully supervised models (see Figure 2), achieving macro F1 scores similar to those of other deep learning approaches for sleep staging [36, 16]. In conclusion, our pretraining method was most beneficial in situations with limited training data, where it outperformed models trained from scratch, but had less impact in situations with large amounts of training data.
We further observed that, while the frequency content of a signal is crucial for sleep staging, deep neural networks extract additional information from the data that exceeds the frequency domain. The importance of the frequency content of a signal for sleep staging is demonstrated by the high macro F1 scores achieved by the fixed feature extractor configuration and the substantial performance improvements it achieved over untrained feature extractors (see Figure 2). We attribute the performance gap between the two training configurations to the priors learned during pretraining the feature extractor. These priors biased the model to extract frequency information from the data, which it achieved with high accuracy after pretraining (see Figure 4b,c). Our finding is consistent with previous studies that reported frequency-based features to be important for sleep staging [33]. When the feature extractor of our model was allowed to be fine-tuned after pretraining, model performance increased (see Figure 2). We hypothesize that this increase in the macro F1 score is due to the feature extractor learning to extract information beyond the frequency content of the signal during fine-tuning. This hypothesis is in line with the AASM annotation guidelines [5], which consider several frequency-unrelated features essential for sleep staging. These features include time-domain information, which is important for spindles as well as amplitudes and specific patterns, such as k-Complexes [15]. Recent studies that applied feature engineering approaches to sleep staging further support our hypothesis by including additional features from the time-domain in the models used [43, 10]. In their work, Vallat and Walker analyzed the most important features for their model and found that time-related features, such as the time elapsed from the beginning of a recording, were among the top 20 most important features [43]. Although it remains unclear what additional information our pretrained models learn during fine-tuning, our method offers a promising avenue for future research into the interpretability of deep neural networks for sleep staging.
There are several opportunities for future work that could build upon our findings. One promising direction is to explore the pretraining task in more detail, for example, by investigating the synthetic data generation process and the impact of changing the used frequency range. Similar to previous work in the vision domain [3], it could also be promising to investigate which structural properties of synthetic time series are important for sleep staging. This could be achieved by defining new pretraining tasks that are based on different data generation processes that incorporate more complex structures like desynchronized phases across channels, noise, or polymorphic amplitude variations. Exploring such data generation processes may lead to a better understanding of what constitutes “natural” EEG time series and what information, besides the frequency content, is essential for sleep staging. In addition, we suggest exploring models with greater capacity and less inductive bias than the CNN-based architecture used in this work, such as transformer models [45, 6], which we expect to benefit even more from our pretraining method. Pretraining such models with synthetic data may alleviate their need for large amounts of training data [11]. Another avenue for future research is to investigate whether our pretraining method is beneficial for specific cohorts of subjects, such as patients with a specific disorder or specific age groups. Although we did investigate datasets with different demographics in this work, we did not perform detailed analyses of the impact of these demographics on model performance. Finally, it could be insightful to compare our approach with a broader range of SSL methods [32] and data augmentation strategies that employ synthetic EEG generators [31, 22]. To enable such comparisons and to facilitate future research in this direction, we make our code available online [19].
Our method presents a novel solution to address important issues that affect current deep learning models in the EEG time series domain, without requiring large amounts of patient data. We expect our approach to be advantageous in various applications where EEG data is scarce or derived from a limited number of subjects, such as brain–computer interfaces [30] or neurological disorder detection [44].
4 Materials and Methods
We trained deep neural networks in two phases: a pretraining phase based on synthetic data and a fine-tuning phase based on clinical sleep staging data. The training process is illustrated in Figure 1.
4.1 Phase 1: Pretraining with Synthetic Data
In the pretraining phase, we trained a deep neural network to predict the frequencies of randomly-generated synthetic time series (“frequency pretraining”) [20]. The deep neural network consisted of a feature extractor to produce features and a classifier to predict which frequencies are present in the synthetic time series. Instead of training the model to predict the exact frequencies, we defined frequency ranges (frequency bins) and let the model identify the bins containing frequencies present in the signal (therefore turning the problem into a multi-label classification task).
4.1.1 Synthetic Data Generation
The synthetic samples used for pretraining consisted of a time series signal $s(t)\in\mathbb{R}^{3\times 3000}$, featuring 3 channels of 30 s sampled at 100 Hz, and an associated label vector $l$. The synthetic time series were created by summing sine waves with randomly sampled frequencies, while the label vector encoded from which frequency ranges (frequency bins) the frequencies were sampled. Since we wanted to test our method in the context of sleep staging, we followed the American Academy of Sleep Medicine (AASM) guidelines and focused on the spectral band between 0.3–35 Hz. This range was divided into 20 frequency bins on a base-2 logarithmic scale, resulting in a label vector $l\in\{0,1\}^{20}$ with binary entries for each bin that indicated whether frequencies in a given bin were present (1) or absent (0) in the time series. To generate a synthetic time series, we first randomly selected a subset of the 20 frequency bins, setting the corresponding entries in the label vector $l$ to 1 (each bin was selected with a probability of 50%). Next, we independently sampled frequencies from the selected bins for each channel and generated sine waves with these frequencies. The sine waves were then shifted by a random phase sampled from a uniform distribution between 0 and $2\pi$, summed, and normalized to create the synthetic signal. Therefore, one channel $s_{c}(t)$ of a synthetic signal $s(t)=(s_{1}(t),s_{2}(t),s_{3}(t))$ is defined as
$s_{c}(t)=\frac{\tilde{s}_{c}(t)-\mu_{c}}{\sigma_{c}},\quad\tilde{s}_{c}(t)=\sum_{i}^{n_{f}}I_{i}\,\sin(2\pi\,tf_{c,i}+\phi_{i}),\quad I_{i}\in\mathcal{U}\{0,1\},$ (1)
where $t$ denotes time (in seconds), $n_{f}$ denotes the number of frequency bins, $I_{i}$ denotes the indicator function encoding whether a frequency within a bin $i$ is present (1) or not (0), $f_{c,i}$ denotes the frequencies of the sine waves (in Hertz), $\phi_{i}$ denotes their phase (in radians), and $\mu_{c}$ and $\sigma_{c}$ denote mean and standard deviation of $\tilde{s}_{c}$. Note that the phases $\phi_{i}$ were sampled for each frequency bin independently but remained the same across channels to keep the data generation process simple, while the frequencies $f_{c,i}$ were sampled independently for each bin and channel.
For each training run, we generated 101,000 synthetic samples, using 100,000 samples for training and the remaining 1000 samples as a validation set to track various metrics. The number of synthetic samples, frequency bins, and the logarithmic scale that defines the boundaries between the bins were determined through preliminary experiments where we explored a wide range of hyperparameters.
4.1.2 Model